### v1

@@ 1 Introduction
The seventh iteration of the ABAW Competition will take place as part of the ABAW Workshop, held in conjunction with ECCV 2024. This competition continues the successful legacy of previous ABAW Competitions, which have been held alongside IEEE CVPR 2024, 2023, 2022 and 2017, ECCV 2022, ICCV 2021 and IEEE FG 2020. These events have consistently attracted numerous teams from academia and industry worldwide [127, 121, 106, 92, 98, 105, 111, 130, 45, 129, 127, 92, 109, 114, 20, 77, 76, 130, 94, 36, 11, 67, 126, 14, 8, 6, 12, 123, 97, 102, 120, 108, 31, 1, 81, 64, 17, 16, 112, 71, 18, 72, 82, 29, 24, 10, 88, 73, 119, 79, 90, 33, 86, 125, 113, 35, 83, 110, 27, 104, 25, 96, 103, 30, 9, 22, 95, 7, 19, 68, 101, 78, 91, 69, 122, 65, 28, 66, 49, 75, 107, 99, 34, 93, 117, 15, 128, 84, 124, 128, 115, 70, 80, 107, 116, 87, 100, 23, 55, 45, 41, 26, 85, 39, 2, 21, 89, 4, 32, 5, 74, 3, 50, 63, 56, 57, 62, 51].

### v2

@@ 1 Introduction
The seventh iteration of the ABAW Competition will take place as part of the ABAW Workshop, held in conjunction with ECCV 2024. This competition continues the successful legacy of previous ABAW Competitions, which have been held alongside IEEE CVPR 2024, 2023, 2022 and 2017, ECCV 2022, ICCV 2021 and IEEE FG 2020. These events have consistently attracted numerous teams from academia and industry worldwide [110, 104, 89, 75, 81, 88, 94, 113, 37, 112, 110, 75, 92, 97, 16, 61, 60, 113, 77, 30, 7, 52, 109, 10, 4, 2, 8, 106, 80, 85, 103, 91, 26, 1, 65, 49, 13, 12, 95, 56, 14, 57, 66, 24, 19, 6, 72, 58, 102, 63, 73, 27, 70, 108, 96, 29, 67, 93, 22, 87, 20, 79, 86, 25, 5, 17, 78, 3, 15, 53, 84, 62, 74, 54, 105, 50, 23, 51, 41, 59, 90, 82, 28, 76, 100, 11, 111, 68, 107, 111, 98, 55, 64, 90, 99, 71, 83, 18, 37, 21, 69].